# Motion blur; Blind deblurring; Blind motion deblurring; Deep learning

### v1

@@ 4 Comparison with State-of-the-Art Methods
As shown in Table 2, this review summarizes over 30 state-of-the-art blind motion deblurring methods based on deep learning. It includes publication details, summaries, advantages, and limitations. Among these, 22 were published in top-tier computer vision conferences such as CVPR, and ICCV, while 8 were featured in leading image processing journals like TIP, and TCSVT, among others.
@@
Table 6: Comparison of generalization performance on real motion blur datasets.
Methods Trained on GoPro Trained on Realblur
RealBlur-J Realblur-R RealBlur-J Realblur-R
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
SRN Tao et al. (2018) 28.56 0.867 35.66 0.947 31.38 0.909 38.65 0.965
SDWNet Zou et al. (2021) 28.61 0.867 35.85 0.948 30.73 0.896 38.21 0.963
DeblurGAN-V2 Kupyn et al. (2019) 28.70 0.866 35.26 0.944 29.69 0.870 36.44 0.935
MPRNet Zamir et al. (2021) 28.70 0.873 35.99 0.952 31.76 0.922 39.31 0.972
MSSNet Kim et al. (2022) 28.79 0.879 35.93 0.953 32.10 0.928 39.76 0.972
LaKDNet Ruan et al. (2023) 28.78 0.878 35.91 0.954 32.33 0.929 39.91 0.974
MAXIM-3S Tu et al. (2022) 28.83 0.875 35.78 0.947 32.84 0.935 39.45 0.962
DeepRFT+ Mao et al. (2023) 28.97 0.884 36.06 0.954 32.19 0.931 39.84 0.972
@@ 5 Current Challenges and Future Prospects
5.1 Challenges of Blind Motion Deblurring and Major Concerns
@@
Despite the progress made in image deblurring algorithms concerning restoration effects, challenges persist due to the complexity of real-world blur scenarios. The challenges of blind motion deblurring and major concerns are as follows:
(1) Generalization Ability of Model: As shown in Table 6, the model trained on synthetic datasets exhibits poor generalization on real datasets. Compared with the training on the real datasets, the overall performance of the methods trained on the synthetic dataset decreases by about 3%, and the SSIM value will decrease by at least 0.02 in Table 6. Enhancing the generalization of model is a question that needs to be addressed.
(2) Data Requirements and Diversity: Most deep learning-based image deblurring algorithms rely on training with clear-blurred image pairs. The types of motion blur are diverse. Handling different types and intensities of motion blur requires a large amount of diverse data. Collecting and labeling this data might be a challenge, especially considering the types of blur, lighting conditions, and scene variations.
@@
(4) Deblurring Methods for High-Level Visual Tasks: Blurred images hinder several advanced computer vision tasks, yet many existing deblurring methods aim solely at recovering high-quality images without considering aiding higher-level tasks such as object detection Zheng et al. (2021); Wu et al. (2023); Zhao et al. (2023) or image segmentation Pan et al. (2016); Luo et al. (2021). Research focusing on effectively extracting relevant features for high-level visual problems to enhance image deblurring solutions is a promising direction.
5.2 Future Prospects
(1) More comprehensive datasets: Most deep learning-based image deblurring algorithms rely on training data comprising blurred images. Constructing a more comprehensive dataset that encompasses a broader range of application scenarios can empower models to better handle various real-world situations, such as motion blur, low-light conditions, and lens shake. However, building such a comprehensive dataset poses challenges. Collecting and labeling various types of blurry images may require substantial time and resources. Moreover, ensuring the quality and diversity of the dataset is crucial, as low-quality or imbalanced data could adversely affect model training.
(2) More innovative network structure: Considering the complexity of image deblurring and the practical demands of applications, it’s essential to innovate and improve the deblurring network architectures. Combining the latest advancements in deep learning, these architectures may better capture complex blur patterns within images. This enhancement could enable them to address various types and intensities of blur, allowing models to perform exceptionally well across a broader range of scenarios.
(3) More objective evaluation indicators: Current evaluation metrics may fail to capture all subtle variations in the image deblurring process. Hence, more objective and comprehensive metrics are crucial for assessing algorithm effectiveness and performance enhancements. Considering multiple aspects such as clarity, detail retention, noise, contrast, etc., constructing comprehensive evaluation metrics would better assess the effectiveness of image deblurring. Introducing evaluation metrics aligned with human visual perception, beyond pixel-level metrics, considering the human perception of details, contrast, color, etc., might better reflect the performance of image deblurring techniques.
(4) Few-shot learning: Traditional deep learning models often require extensive labeled data for training, which might be challenging to obtain, especially high-quality labeled clear images. Few-shot learning Zhang et al. (2019); Chi et al. (2021); Yang et al. (2022) help models rapidly learn and adapt to new tasks from a limited amount of labeled data, reducing dependence on extensive labeled datasets. This approach addresses challenges related to acquiring and labeling data, potentially widening the application of image deblurring techniques across real-world scenarios.
(5) Applications of diffusion: The Diffusion model is a generative model that has garnered significant attention in recent years, demonstrating immense potential in the realm of image editing and enhancement. In the field of image deblurring, the Diffusion model has also found applications  Ren et al. (2023); Whang et al. (2022). Its flexibility allows it to handle various types of noise and transformations, suggesting its capability to potentially address complex image blurring. In the future, with a deeper understanding of the Diffusion model and the development of related techniques, it is poised to become a significant tool in the domain of image deblurring, promising more satisfying results and improvements.
(6) Unsupervised Deblurring Methods: Real-world blurry images often lack corresponding clear counterparts. Employing unsupervised methods Zhao et al. (2022) could render algorithms robust against real-world blurry scenarios, operating without extensive training data.
@@ 6 Conclusions
The current research and developments indicate significant progress in enhancing image quality and clarity through deep learning technologies. This paper reviews recent advancements in deep learning-based image deblurring algorithms, providing a comprehensive classification and analysis. It elaborates on existing image deblurring algorithms from four perspectives: CNN, RNN, GAN, and Transformer. Additionally, it conducts comparative analyses of representative algorithms, summarizing their strengths and limitations. Finally, the paper outlines future research directions in image deblurring algorithms. However, it is acknowledged that addressing more complex blurs in real-world scenarios remains a challenge. The quality of data and the accuracy of labels are crucial for training deep learning models, necessitating large-scale, high-quality datasets to support model training and optimization. In the future, there is an anticipation to further optimize deep learning models to enhance their speed and efficiency, expanding the application of these techniques in fields such as video processing, surveillance, and autonomous driving.
[…]
Methodology, Y.X.; Supervision, Z.L. and Y.X.; Writing—Original Draft, Y.X., H.Z. and C.L.; Writing—Review and Editing, Y.X., H.Z. and F.S.; Visualization, Y.X., F.S., C.L. and H.Z.; All authors have read and agreed to the published version of the manuscript.
[…]
This research received no external funding.
[…]
Not applicable.
[…]
The authors declare no conflict of interest.

### v2

@@ 4 Comparison with State-of-the-Art Methods
Table 5: Speed and Parameters of representative SOTA methods.
Method Category Param. (M) Time (s) PSNR (%)
DMPHN [87] CNN 21.7 0.21 31.20
MIMOU-Net+ [85] CNN 16.1 0.02 32.45
MPRNet [90] CNN 23.0 0.09 32.66
SRN [81] RNN 6.8 0.07 30.26
DeblurGAN-V2 [103] GAN 60.9 0.04 29.55
Restormer [125] Transformer 26.1 0.08 32.92
Uformer [121] Transformer 50.9 0.07 32.97
Stripformer [124] Transformer 19.7 0.04 33.08
FFTformer [126] Transformer 16.6 0.13 34.21
@@
As shown in Table LABEL:tab:advantages, this review summarizes over 30 state-of-the-art blind motion deblurring methods based on deep learning. It includes publication details, summaries, advantages, and limitations. Among these, 22 were published in top-tier computer vision conferences such as CVPR, and ICCV, while 8 were featured in leading image processing journals like TIP, and TCSVT, among others.
@@ 5 Current Challenges and Future Prospects
Despite the progress made in image deblurring algorithms concerning restoration effects, challenges persist due to the complexity of real-world blur scenarios. The challenges of blind motion deblurring and future prospects are as follows:
@@
(2) Deblurring Methods for High-Level Visual Tasks: Blurred images hinder several advanced computer vision tasks, yet many existing deblurring methods aim solely at recovering high-quality images without considering aiding higher-level tasks such as object detection [163, 164, 165] or image segmentation [166, 167]. Research focusing on effectively extracting relevant features for high-level visual problems to enhance image deblurring solutions is a promising direction.
(3) Few-shot Learning: Traditional deep learning models often require extensive labeled data for training, which might be challenging to obtain, especially high-quality labeled clear images. Few-shot learning [168, 169, 169] help models rapidly learn and adapt to new tasks from a limited amount of labeled data, reducing dependence on extensive labeled datasets. This approach addresses challenges related to acquiring and labeling data, potentially widening the application of image deblurring techniques across real-world scenarios.
(4) Applications of Diffusion: The Diffusion model is a generative model that has garnered significant attention in recent years, demonstrating immense potential in the realm of image editing and enhancement. In the field of image deblurring, the Diffusion model has also found applications [170, 171]. Its flexibility allows it to handle various types of noise and transformations, suggesting its capability to potentially address complex image blurring. In the future, with a deeper understanding of the Diffusion model and the development of related techniques, it is poised to become a significant tool in the domain of image deblurring, promising more satisfying results and improvements.
@@ 6 Conclusions
The current research and developments indicate significant progress in enhancing image quality and clarity through deep learning technologies. This paper reviews recent advancements in deep learning-based image deblurring algorithms, providing a comprehensive classification and analysis. It elaborates on existing image deblurring algorithms from four perspectives: CNN, RNN, GAN, and Transformer. Additionally, it conducts comparative analyses of representative algorithms, summarizing their strengths and limitations. Finally, the paper outlines future research directions in image deblurring algorithms. However, it is acknowledged that addressing more complex blurs in real-world scenarios remains a challenge. The quality of data and the accuracy of labels are crucial for training deep learning models, necessitating large-scale, high-quality datasets to support model training and optimization. In the future, there is an anticipation to further optimize deep learning models to enhance their speed and efficiency, expanding the application of these techniques in fields such as video processing, surveillance, and autonomous driving.All collected models, benchmark datasets, source code links, and codes for evaluation have been made publicly available at https://github.com/VisionVerse/Blind-Motion-Deblurring-Survey.
Declarations
  • Funding:This research received no external funding.
  • Author contribution: Methodology, Y.X.; Supervision, Z.L. and Y.X.; Writing—Original Draft, Y.X., H.Z. and C.L.; Writing—Review and Editing, Y.X., H.Z. and F.S.; Visualization, Y.X., F.S., C.L. and H.Z.; All authors have read and agreed to the published version of the manuscript.